OpenStudy (anonymous):

1/4x-5=x please help me solve this

OpenStudy (harsimran_hs4):

can you please use equation generator or use brackets so that question is clear

OpenStudy (anonymous):

x = -20/3 I believe @meaah_c_(:

OpenStudy (anonymous):

(1/4x)-5=x

OpenStudy (anonymous):

@meaah_c_:

OpenStudy (harsimran_hs4):

NOTE : x can never take value 0 as this will make the expression not defined rearranging a bit (1/4x) = x + 5 multiplying by 4x on both sides 1 = 4x(x + 5) can you proceed now ?

OpenStudy (anonymous):

Thank you glad I could help. :)